# A Systematic Survey on Large Language Models for Evolutionary Optimization: From Modeling to Solving

### v2

@@ IV-A Prompt-based Methods
Automated optimization modeling, which aims to formalize real-world challenges into mathematical models, has recently emerged as a significant research direction. To mitigate the reliance on expert experience during this process, researchers have explored various approaches. For instance, Pawlak et al. preliminarily demonstrated the feasibility of automated modeling by analyzing feasible solution samples [97] or employing evolutionary algorithms [98] to generate constraints. However, these methods were often constrained by sample quality and limited in scope [69], primarily focusing on constraint generation rather than a direct conversion from natural language descriptions to complete mathematical models.
@@
To address this research gap, Ramamonjison et al. proposed OptGen [99], the first method to convert unstructured natural language problem descriptions into mathematical optimization models, and subsequently catalyzed the field’s growth with the NL4OPT competition [100]. This competition decomposed the task into two core stages: first, using Named Entity Recognition (NER) to extract key entities such as variables, parameters, and constraints from the problem text, and then generating a complete mathematical model from the entity-annotated description. However, these earlier works largely relied on lightweight pre-trained models like BERT [101, 102] and BART [103, 104]. In contrast, with the advent of LLMs, B. Almonacid et al. were the first to attempt automated modeling using powerful LLMs such as GPT-3.5 [67]. They employed a single-step modeling approach, directly prompting the LLM to output the mathematical model corresponding to a problem description and verifying its correctness with a solver. Nevertheless, while this single-step method proves effective for simple problems, its accuracy and robustness are generally insufficient for meeting the high standards required for complex optimization modeling.
@@
Inspired by the aforementioned studies, researchers have progressively adopted large language models for a two-stage approach involving NER and automated modeling in the NL4OPT task. The Holy Grail 2.0 framework [68] proposed by Tsouros et al., for instance, recommends using specialized models like Ner4OPT for the NER stage. It then further decomposes the automated modeling task into three sub-tasks: first, identifying relationships between entities; second, formalizing the problem into an optimization model using these entities and their relationships; and finally, generating code for testing and execution. Building upon this idea, Li et al. [69] proposed a framework specifically for Mixed-Integer Linear Programming (MILP). In their approach, they employed a fine-tuned model to classify constraints to handle logical constraints and binary variables, which were less explored in previous works. Furthermore, this two-stage method has been successfully applied in various specialized domains. For example, [73] applied it to the field of energy management, requiring the LLM to first extract the five most critical parameters from the problem description before proceeding with automated modeling. Another study [77] applied this approach to travel planning, where the LLM was responsible for extracting user preferences and associated costs, building a complete model, and utilizing a solver to find the optimal travel plan.
Beyond the two-stage framework, researchers have actively explored other prompt-based paradigms to further enhance the performance of large language models in optimization modeling. Two of the most representative paradigms are the multi-agent framework and the interactive framework. The former decomposes the automated modeling task into a series of collaborative expert sub-tasks, which are then completed jointly by multiple LLMs. The latter, on the other hand, aims to alleviate the user’s input burden by not requiring a complete problem description upfront. Instead, the optimization model is progressively refined through human-computer interaction, allowing for incremental additions or removals of constraints, priority adjustments, and other modifications.
@@
In the domain of multi-agent frameworks, the Chain-of-Experts [70] framework proposed by Xiao et al. is a foundational work. They initialized 11 types of experts, including a terminology expert, a modeling expert, a programming expert, and a code review expert, which were coordinated by a conductor agent. This framework dynamically and iteratively selects experts to construct a forward chain-of-thought and incorporates a backward reflection mechanism to handle errors and inconsistencies during the modeling process. Building upon this, subsequent studies have further explored diverse multi-agent collaboration modes [71, 79, 85, 86]. For instance, AhmadiTeshnizi et al. introduced OptiMUS [71] and OptiMUS-0.3 [79], which also use a conductor agent to coordinate modeling, programming, and evaluation tasks. However, this manager reviews the content of each conversational step before assigning the corresponding task to the next agent. However, Wan et al. pointed out that this conductor-orchestrated framework could lead to issues such as an unpredictable workflow and insufficient mathematical accuracy. Therefore, they proposed a cognitively-inspired framework named ORMind [85], which aims to improve the reliability and transparency of solutions while reducing API calls by mimicking the human cognitive process and adopting a structured, predictable workflow alongside a counterfactual reasoning mechanism. Additionally, Mostajabdaveh et al. [72] focused more on the evaluation phase of modeling, replacing the time-consuming process of traditional solver-based verification with peer verification among agents. Similarly, Talebi et al. [84] also centered their research on the evaluation phase, specifically for stochastic optimization problems, by setting up multiple independent reviewer agents to score the modeling results, thereby aiming for a comprehensive assessment of modeling quality.
@@
Both the two-stage and multi-agent frameworks require users to provide a complete problem description before the framework is initiated. However, for many real-world optimization problems, such as conference scheduling and travel planning, user preferences can change dynamically, which necessitates a paradigm that supports continuous interaction between the user and the model. To address this, Zhang et al. proposed the OptLLM [78] framework, which supports both single-shot and interactive inputs. This framework uses an interactive refinement module to help users progressively refine the problem description, while a converter module and a response module handle the subsequent modeling and verification tasks. Lawless et al. [80] further advanced this interactive framework by requiring the LLM to choose from five tasks, such as adding or removing constraints and adjusting priorities, after each user input. Their MeetMeta system not only provides a solution but also explains why certain preferences cannot be met and allows users to make modifications, creating a continuous feedback loop. Subsequent research [81] has focused more on how to translate user priorities into optimization constraints during a dialogue, thereby enabling users to explore trade-offs in tasks like ridesharing coordination.
@@
As research into prompt-based methods has deepened, researchers have begun exploring different avenues to address various existing challenges in LLM-based automated modeling. On the modeling generation side, to tackle the vast and problem-dependent hypothesis space faced by LLMs, researchers have adopted distinct strategies. For instance, Astorga et al. [82] combined LLMs with Monte Carlo Tree Search (MCTS) to efficiently explore this hierarchical and immense hypothesis space, further enhancing search efficiency through symbolic pruning and LLM-based evaluation. In contrast to this search-first strategy, Li et al. [76] employed a generation-first approach. They constrained the LLM’s output using predefined structures, thereby directly avoiding blind searches within the huge hypothesis space. This was complemented by a post-processing verification and correction mechanism to ensure the generated models could be correctly parsed by solvers. On the model verification side, to improve the effectiveness and broaden the scope of validation, Huang et al. [75] proposed an automated evaluation framework based on solver outputs. They also constructed the MAMO benchmark, which extends the NL4OPT benchmark by including systems of ordinary differential equations as well as more complex LP and MILP problems. Conversely, Wang et [105] al. took a different approach, employing a modified Weisfeiler-Lehman (WL) graph isomorphism algorithm to verify model equivalence by transforming optimization models into a bipartite graph structure. They also created the OptiBench benchmark, which contains a large number of LP and MILP problems designed to effectively reflect the performance boundaries of LLMs. Similarly, Zhai et al. [83] introduced the Quasi-Karp equivalence theoretical framework and developed the EquivaMap system, achieving a scalable method for automated equivalence detection through LLM-generated variable mapping functions and lightweight verification.
@@
Despite the significant potential demonstrated by prompt-based frameworks in automated optimization modeling, their technical boundaries remain constrained by the core capabilities of the underlying LLMs. Current mainstream approaches, which rely on closed-source models like the GPT series, face a dual challenge [27]: on one hand, the scarcity of high-quality training data in the optimization modeling domain severely restricts the models’ performance ceiling; on the other hand, using API calls to commercial models may lead to the leakage of sensitive information, posing security risks in high-stakes scenarios such as medical scheduling and defense planning. These fundamental limitations, coupled with the unique formal rigor required for optimization problems, are driving a paradigm shift in research from prompt engineering to learning-based methods. The latter, through fine-tuning open-source models, can achieve more precise mathematical expression generation, more controllable data privacy protection, and more efficient injection of domain knowledge.
@@ IV-B Learning-based Methods
Learning-based methods fundamentally overcome the inherent limitations of prompt engineering by fine-tuning model parameters to internalize optimization modeling knowledge within large language models. In early research, several studies attempted to fine-tune models to enhance specific aspects of automated modeling. For instance, the work [66] combined traditional NLP methods with LLMs to obtain the Ner4OPT model, which is specialized for the NER task. Additionally, Li et al. [69] utilized a fine-tuned model to classify constraints to uncover complex relationships between entities. While these works involved model fine-tuning, their core idea was to serve as auxiliary components within prompt-based two-stage frameworks. Subsequently, the LM4OPT [87] framework proposed by Ahmed et al. significantly improved the optimization modeling capabilities of Llama-2-7b [106] through progressive fine-tuning. However, due to its fine-tuning data being confined to the NL4OPT dataset, its modeling performance still exhibited a substantial gap compared to closed-source models like GPT-4 [107].
@@
Building upon these studies, Huang et al. proposed ORLM [27], a framework featuring a two-stage data synthesis strategy comprising expansion and augmentation. This strategy systematically prepares structured data for model fine-tuning, encompassing three critical components: natural language problem descriptions, formal mathematical models, and executable solver code. Specifically, the expansion stage leverages a small set of seed data and uses GPT-4 to iteratively generate new samples with diverse scenarios and difficulties. The augmentation stage, in turn, further enhances sample variety by employing strategies such as modifying objective functions, altering constraints, and rephrasing problems. During the iterative data synthesis process, the framework also employs regular match correction function to address minor grammatical errors in the code and removes duplicate problems to ensure the integrity and purity of the dataset. Subsequently, they fine-tuned a series of open-source models, such as Mistral-7B [108] and Deepseek-Math-7B-Base [109], using Instruction-Tuning. Comparative tests revealed that the fine-tuned open-source models not only consistently outperformed single-step GPT-4 modeling but also yielded significant performance improvements over classic prompt-based methods like Chain-of-Experts [70] and OptiMUS [71]. The data synthesis-model fine-tuning-evaluation workflow has been widely adopted by subsequent research, with some studies focusing on proposing new data synthesis methods, while others prioritize optimizing the fine-tuning stage.
@@
A subset of learning-based research focuses on the data synthesis component, aiming to address the challenge of data scarcity in optimization modeling. The ReSocratic framework [88] developed by Yang et al. aims to synthesize data in an inverse manner, first synthesizing optimization examples with mathematical formulas and then translating them in reverse into natural language problems, thereby creating a large-scale dataset. Similarly, the OptMATH framework [91] proposed by Lu et al. is a scalable bidirectional data synthesis framework that automatically generates problems with controllable complexity from seed data and uses a forward modeling and rejection sampling mechanism to verify the accuracy of the data pairs. Zhou et al. combined the diversity of forward generation and the reliability of inverse generation from the aforementioned studies to design the DualReflect synthetic data pipeline [95], which is specifically tailored for dynamic programming problems. Beyond focusing on the synthesis methods themselves, other researchers have conducted in-depth studies on the quality of the synthesized content. Wang et al. argued that the data paradigm of methods like ORLM has too large a span from natural language to mathematical models. To address this, they constructed the StructuredOR dataset [90], which aims to solve the problem of missing detailed modeling process annotations (such as variable definitions) in existing datasets, providing more comprehensive annotation information for model learning. Wu et al. focused on generating more complex optimization problems. Their Step-Opt framework [94] systematically increases problem complexity through an iterative problem generation approach and employs a stepwise validation mechanism to ensure data rigor.
@@
Apart from data synthesis, another line of research focuses on fine-tuning methods. Jiang et al. [89] enhanced the model’s accuracy and generalization in problem formalization and solver code generation through multi-instruction fine-tuning, while also introducing model alignment and self-correction mechanisms to mitigate hallucination phenomena. Chen et al. proposed the SIRL framework [93] by incorporating reinforcement learning, which utilizes an external optimization solver as a validator to provide verifiable reward signals for the reinforcement learning process. This approach effectively addresses the hallucination problem in LLM generation, thereby improving the factual correctness and trustworthiness of the generated content. Furthermore, Amarasinghe et al. [92] proposed a method for fine-tuning cost-effective LLMs to tackle specific business challenges such as production scheduling, which demonstrated superior performance in real-world applications.

### v3

@@ IV-A Prompt-based Methods
Automated optimization modeling, which aims to formalize real-world challenges into mathematical models, has recently emerged as a significant research direction. To mitigate the reliance on expert experience during this process, researchers have explored various approaches. For instance, Pawlak et al. preliminarily demonstrated the feasibility of automated modeling by analyzing feasible solution samples [pawlak2017automatic] or employing evolutionary algorithms [pawlak2021grammatical] to generate constraints. However, these methods were often constrained by sample quality and limited in scope [li2023synthesizing], primarily focusing on constraint generation rather than a direct conversion from natural language descriptions to complete mathematical models.
@@
To address this research gap, Ramamonjison et al. proposed OptGen [ramamonjison2022augmenting], the first method to convert unstructured natural language problem descriptions into mathematical optimization models, and subsequently catalyzed the field’s growth with the NL4OPT competition [ramamonjison2023nl4opt]. This competition decomposed the task into two core stages: first, using Named Entity Recognition (NER) to extract key entities such as variables, parameters, and constraints from the problem text, and then generating a complete mathematical model from the entity-annotated description. However, these earlier works largely relied on lightweight pre-trained models like BERT [wang2023opd, doan2022vtcc] and BART [gangwar2023highlighting, ning2023novel]. In contrast, with the advent of LLMs, B. Almonacid et al. were the first to attempt automated modeling using powerful LLMs such as GPT-3.5 [almonacid2023towards]. They employed a single-step modeling approach, directly prompting the LLM to output the mathematical model corresponding to a problem description and verifying its correctness with a solver. Nevertheless, while this single-step method proves effective for simple problems, its accuracy and robustness are generally insufficient for meeting the high standards required for complex optimization modeling.
@@
Inspired by the aforementioned studies, researchers have progressively adopted large language models for a two-stage approach involving NER and automated modeling in the NL4OPT task. The Holy Grail 2.0 framework [tsouros2023holy] proposed by Tsouros et al., for instance, recommends using specialized models like Ner4OPT for the NER stage. It then further decomposes the automated modeling task into three sub-tasks: first, identifying relationships between entities; second, formalizing the problem into an optimization model using these entities and their relationships; and finally, generating code for testing and execution. Building upon this idea, Li et al. [li2023synthesizing] proposed a framework specifically for Mixed-Integer Linear Programming (MILP). In their approach, they employed a fine-tuned model to classify constraints to handle logical constraints and binary variables, which were less explored in previous works. Furthermore, this two-stage method has been successfully applied in various specialized domains. For example, [jin2024democratizing] applied it to the field of energy management, requiring the LLM to first extract the five most critical parameters from the problem description before proceeding with automated modeling. Another study [de2024trip] applied this approach to travel planning, where the LLM was responsible for extracting user preferences and associated costs, building a complete model, and utilizing a solver to find the optimal travel plan.
Beyond the two-stage framework, multi-agent frameworks extend the idea by replacing a fixed NER-to-modeling pipeline with a richer set of expert roles coordinated across multiple LLMs, broadening coverage and strengthening cross-checking. Interactive frameworks differ in focus: they turn modeling into an ongoing dialogue with the user, refining the formulation through iterative feedback, constraint insertion or deletion, priority adjustments, and preference updates, so the model evolves dynamically instead of being fixed upfront.
@@
In the domain of multi-agent frameworks, the Chain-of-Experts [xiao2023chain] framework proposed by Xiao et al. is a foundational work. They initialized 11 types of experts, including a terminology expert, a modeling expert, a programming expert, and a code review expert, which were coordinated by a conductor agent. This framework dynamically and iteratively selects experts to construct a forward chain-of-thought and incorporates a backward reflection mechanism to handle errors and inconsistencies during the modeling process. Building upon this, subsequent studies have further explored diverse multi-agent collaboration modes [ahmaditeshnizi2024optimus, ahmaditeshnizi2024optimus03, wang2025ormind, liang2025llm]. For instance, AhmadiTeshnizi et al. introduced OptiMUS [ahmaditeshnizi2024optimus] and OptiMUS-0.3 [ahmaditeshnizi2024optimus03], which also use a conductor agent to coordinate modeling, programming, and evaluation tasks. However, this manager reviews the content of each conversational step before assigning the corresponding task to the next agent. However, Wan et al. pointed out that this conductor-orchestrated framework could lead to issues such as an unpredictable workflow and insufficient mathematical accuracy. Therefore, they proposed a cognitively-inspired framework named ORMind [wang2025ormind], which aims to improve the reliability and transparency of solutions while reducing API calls by mimicking the human cognitive process and adopting a structured, predictable workflow alongside a counterfactual reasoning mechanism. Additionally, Mostajabdaveh et al. [mostajabdaveh2024optimization] focused more on the evaluation phase of modeling, replacing the time-consuming process of traditional solver-based verification with peer verification among agents. Similarly, Talebi et al. [talebi2025large] also centered their research on the evaluation phase, specifically for stochastic optimization problems, by setting up multiple independent reviewer agents to score the modeling results, thereby aiming for a comprehensive assessment of modeling quality.
@@
Both the two-stage and multi-agent frameworks require users to provide a complete problem description before the framework is initiated. However, for many real-world optimization problems, such as conference scheduling and travel planning, user preferences can change dynamically, which necessitates a paradigm that supports continuous interaction between the user and the model. To address this, Zhang et al. proposed the OptLLM [zhang2024solving] framework, which supports both single-shot and interactive inputs. This framework uses an interactive refinement module to help users progressively refine the problem description, while a converter module and a response module handle the subsequent modeling and verification tasks. Lawless et al. [lawless2024want] further advanced this interactive framework by requiring the LLM to choose from five tasks, such as adding or removing constraints and adjusting priorities, after each user input. Their MeetMeta system not only provides a solution but also explains why certain preferences cannot be met and allows users to make modifications, creating a continuous feedback loop. Subsequent research [gomez2025values] has focused more on how to translate user priorities into optimization constraints during a dialogue, thereby enabling users to explore trade-offs in tasks like ridesharing coordination.
@@
As research into prompt-based methods has deepened, researchers have begun exploring different avenues to address various existing challenges in LLM-based automated modeling [deng2024cafa]. On the modeling generation side, to tackle the vast and problem-dependent hypothesis space faced by LLMs, researchers have adopted distinct strategies. For instance, Astorga et al. [astorga2024autoformulation] combined LLMs with Monte Carlo Tree Search (MCTS) to efficiently explore this hierarchical and immense hypothesis space, further enhancing search efficiency through symbolic pruning and LLM-based evaluation. In contrast to this search-first strategy, Li et al. [li2025abstract] employed a generation-first approach. They constrained the LLM’s output using predefined structures, thereby directly avoiding blind searches within the huge hypothesis space. This was complemented by a post-processing verification and correction mechanism to ensure the generated models could be correctly parsed by solvers. On the model verification side, to improve the effectiveness and broaden the scope of validation, Huang et al. [huang2024llms] proposed an automated evaluation framework based on solver outputs. They also constructed the MAMO benchmark, which extends the NL4OPT benchmark by including systems of ordinary differential equations as well as more complex LP and MILP problems. Conversely, Wang et al. [wang2024optibench]. took a different approach, employing a modified Weisfeiler-Lehman (WL) graph isomorphism algorithm to verify model equivalence by transforming optimization models into a bipartite graph structure. They also created the OptiBench benchmark, which contains a large number of LP and MILP problems designed to effectively reflect the performance boundaries of LLMs. Similarly, Zhai et al. [zhai2025equivamap] introduced the Quasi-Karp equivalence theoretical framework and developed the EquivaMap system, achieving a scalable method for automated equivalence detection through LLM-generated variable mapping functions and lightweight verification.
@@
Despite the significant potential demonstrated by prompt-based frameworks in automated optimization modeling, their technical boundaries remain constrained by the core capabilities of the underlying LLMs. Current mainstream approaches, which rely on closed-source models like the GPT series, face a dual challenge [huang2025orlm]: on one hand, the scarcity of high-quality training data in the optimization modeling domain severely restricts the models’ performance ceiling; on the other hand, using API calls to commercial models may lead to the leakage of sensitive information, posing security risks in high-stakes scenarios such as medical scheduling and defense planning. These fundamental limitations, coupled with the unique formal rigor required for optimization problems, are driving a paradigm shift in research from prompt engineering to learning-based methods. The latter, through fine-tuning open-source models, can achieve more precise mathematical expression generation, more controllable data privacy protection, and more efficient injection of domain knowledge.
@@ IV-B Learning-based Methods
Learning-based methods fundamentally overcome the inherent limitations of prompt engineering by fine-tuning model parameters to internalize optimization modeling knowledge within large language models. In early research, several studies attempted to fine-tune models to enhance specific aspects of automated modeling. For instance, the work [dakle2023ner4opt] combined traditional NLP methods with LLMs to obtain the Ner4OPT model, which is specialized for the NER task. Additionally, Li et al. [li2023synthesizing] utilized a fine-tuned model to classify constraints to uncover complex relationships between entities. While these works involved model fine-tuning, their core idea was to serve as auxiliary components within prompt-based two-stage frameworks. Subsequently, the LM4OPT [ahmed2024lm4opt] framework proposed by Ahmed et al. significantly improved the optimization modeling capabilities of Llama-2-7b [touvron2023llama] through progressive fine-tuning. However, due to its fine-tuning data being confined to the NL4OPT dataset, its modeling performance still exhibited a substantial gap compared to closed-source models like GPT-4 [openai2023gpt4].
@@
Building upon these studies, Huang et al. proposed ORLM [huang2025orlm], a framework featuring a two-stage data synthesis strategy comprising expansion and augmentation. This strategy systematically prepares structured data for model fine-tuning, encompassing three critical components: natural language problem descriptions, formal mathematical models, and executable solver code. Specifically, the expansion stage leverages a small set of seed data and uses GPT-4 to iteratively generate new samples with diverse scenarios and difficulties. The augmentation stage, in turn, further enhances sample variety by employing strategies such as modifying objective functions, altering constraints, and rephrasing problems. During the iterative data synthesis process, the framework also employs regular match correction function to address minor grammatical errors in the code and removes duplicate problems to ensure the integrity and purity of the dataset. Subsequently, they fine-tuned a series of open-source models, such as Mistral-7B [DBLP:journals/corr/abs-2310-06825] and Deepseek-Math-7B-Base [shao2024deepseekmath], using Instruction-Tuning. Comparative tests revealed that the fine-tuned open-source models not only consistently outperformed single-step GPT-4 modeling but also yielded significant performance improvements over classic prompt-based methods like Chain-of-Experts [xiao2023chain] and OptiMUS [ahmaditeshnizi2024optimus]. The data synthesis-model fine-tuning-evaluation workflow has been widely adopted by subsequent research, with some studies focusing on proposing new data synthesis methods, while others prioritize optimizing the fine-tuning stage.
@@
A subset of learning-based research focuses on the data synthesis component [lima2025toward], aiming to address the challenge of data scarcity in optimization modeling. The ReSocratic framework [yang2024optibench] developed by Yang et al. aims to synthesize data in an inverse manner, first synthesizing optimization examples with mathematical formulas and then translating them in reverse into natural language problems, thereby creating a large-scale dataset. Similarly, the OptMATH framework [lu2025optmath] proposed by Lu et al. is a scalable bidirectional data synthesis framework that automatically generates problems with controllable complexity from seed data and uses a forward modeling and rejection sampling mechanism to verify the accuracy of the data pairs. Zhou et al. combined the diversity of forward generation and the reliability of inverse generation from the aforementioned studies to design the DualReflect synthetic data pipeline [zhou2025auto], which is specifically tailored for dynamic programming problems. Beyond focusing on the synthesis methods themselves, other researchers have conducted in-depth studies on the quality of the synthesized content. Wang et al. argued that the data paradigm of methods like ORLM has too large a span from natural language to mathematical models. To address this, they constructed the StructuredOR dataset [wang2024bpp], which aims to solve the problem of missing detailed modeling process annotations (such as variable definitions) in existing datasets, providing more comprehensive annotation information for model learning. Wu et al. focused on generating more complex optimization problems. Their Step-Opt framework [wu2025step] systematically increases problem complexity through an iterative problem generation approach and employs a stepwise validation mechanism to ensure data rigor.
@@
Apart from data synthesis, another line of research focuses on fine-tuning methods [jiang2024llmopt, chen2025solver, amarasinghe2023ai, zhou2025steporlm]. Jiang et al. [jiang2024llmopt] enhanced the model’s accuracy and generalization in problem formalization and solver code generation through multi-instruction fine-tuning, while also introducing model alignment and self-correction mechanisms to mitigate hallucination phenomena. Chen et al. proposed the SIRL framework [chen2025solver] by incorporating reinforcement learning, which utilizes an external optimization solver as a validator to provide verifiable reward signals for the reinforcement learning process. This approach effectively addresses the hallucination problem in LLM generation, thereby improving the factual correctness and trustworthiness of the generated content. Furthermore, Amarasinghe et al. [amarasinghe2023ai] proposed a method for fine-tuning cost-effective LLMs to tackle specific business challenges such as production scheduling, which demonstrated superior performance in real-world applications.